# Analyzing PDFs like Binaries: Adversarially Robust PDF Malware Analysis via Intermediate Representation and Language Model

### v1

@@ 6.6. Ablation Study
Impact of Embedding Raw Content on Classification In this experiment, we first pre-trained the raw content of the PDF object, shown in Figure 1(b), using the standard BERT pre-training method with default preprocessing and tokenization configurations. This pre-training was performed on the baseline dataset (with a corpus size of 20k samples). Then, we applied the general embedding schemes from §$\S$5.3 to embed the raw content of PDF objects, obtaining node embeddings. We then evaluated their performance on the PDF malware classification task, maintaining the same GIN model and training settings as described in §$\S$5.4. The results, presented in the first three rows of Table 6, indicate that the BERTR scheme, pre-trained on raw content, performs worse than the BERT scheme pre-trained on PDFObj IR (as shown in Table 4). Specifically, with a corpus size of 20k, the accuracy on the baseline dataset decreased by 0.42%, and on the extended dataset, it dropped by 2.07%. Furthermore, the results show that directly applying general embedding schemes to raw content performs worse than applying them at the PDFObj IR level. For instance, the BERT BaseR scheme demonstrated a significant performance drop: compared to BERT Base in Table 4, accuracy on the baseline and extended datasets decreased by 1.29% and 8.56%, respectively. Similarly, CodeT5R saw accuracy reductions of 0.18% and 2.84%, respectively. Detailed results are provided in Appendix I.
@@
Impact of ORG Structure on Classification To evaluate the impact of graph structures on classifier robustness, we removed the ORG structure and GIN classifier. Given the variability in the number of objects across PDF samples, we represented each PDF by averaging the semantic vectors of all objects and used this representation as input to a DNN classifier. We selected four PDFObj2Vec schemes that demonstrated some adversarial robustness, as discussed in §$\S$6.5. We trained the DNN classifier on the baseline dataset and assessed its performance. The experimental results highlight the critical importance of the ORG structure in enhancing both the performance and robustness of PDF malware classification. When the ORG structure was removed and a DNN classifier was used with semantic vectors of PDF objects, the classification accuracy on both the baseline and extended datasets declined. Furthermore, the DNN classifiers completely failed to exhibit resistance to adversarial samples, rendering them unable to detect such attacks effectively.
@@ 7. Discussion & Conclusion
Obfuscation Obfuscation is a common tactic in malware to evade detection, and similar techniques, such as keyword obfuscation, are used in PDF malware to bypass detection mechanisms (Stokes, 2019; Lu et al., 2013; Pires, 2020). For example, replacing Name /URI with /#55RI, prevents classifiers that rely on keyword statistical features from obtaining correct statistical features and also hinders analyzers that use Hidost path features from extracting the correct path features. Our method mitigates such obfuscation by retaining the obfuscated keyword data in the PDFObj IR corpus, allowing us to learn the obfuscated semantics. Additionally, we integrate the structural features of ORG to classify PDF malware, enhancing the precise detection of obfuscated PDF malware.
Concept Drift Over time, the effectiveness of traditional machine learning models trained on outdated datasets tends to decline (Jordaney et al., 2017; Yang et al., 2021; Barbero et al., 2022). As multimedia evolves, benign PDFs are advancing faster than malicious ones, altering their characteristics significantly. Consequently, SOTA classifiers, including RTC (Chen et al., 2020) and CFTC (Tong et al., 2019), struggle to accurately identify new benign samples of the extended dataset, leading to lower accuracy, as shown in Table 4. Since PDFObj2Vec is trained on PDFObj IR, a fundamental representation of PDF objects, the basic changes in PDFObj IR are minimal, regardless of variations in the PDF structure. Therefore, our proposed method is minimally affected by concept drift. The experimental results presented in Table 4 demonstrate that our proposed approaches, which leverage PDFObj IR, not only outperform methods that do not utilize this representation but also surpass SOTA classifiers across all metrics on the most recent extended dataset.
@@ B. Error Handling and Correction
Apart from the three main errors discussed in §$\S$4.3, there are other errors that we have identified. These additional errors and their handling measures are listed in Table 8.
@@ C. Details of Pre-trained PDFObj2Vec
Word2Vec-Based Scheme For a sentence structure like Type_name Parent_ref MediaBox_num_list Resources_dict Resources/P-rocSet_name_list Resources/XObject_DICT... , suppose we select MediaBox_num_list as the target word, and the context window size is 2. This means we will use the two words before and after this word as the context, as shown in Figure 11. Thus, the context words are: [Type_name Parent_ref Resources_dict Resources/ProcSet_name_list]. The overall semantics of the context are represented by calculating the average of the context word vectors. This context vector is then fed into a softmax classifier, which predicts the probability of each word being the target word. In §$\S$6, our implementation of Word2Vec consists of two embedding layers, one for learning the embeddings of the center words and the other for the context words, along with a Dropout layer. We update the word vectors using the cross-entropy loss between the target words and the model’s predictions. The optimizer used is Stochastic Gradient Descent (SGD), with a learning rate of 5e-4.
@@ E. Details of GIN Classifier
We trained our implemented GIN classifier on the baseline dataset using the Deep Graph Library (Library, 2024). The classifier consists of two GIN layers, each with a hidden layer size of 256, employing mean aggregation and the ReLU activation function. The update rule for a GIN graph convolution layer can be expressed as:
@@ H. Additional Details of Adversarial Attack
We summarize the adversarial attacks we use in §$\S$6.5, as shown in Table 11. Specifically, the gradient-based attack, genetic algorithm-based attack, and random noise attack are conducted in the feature space, where attackers are assumed to have varying degrees of access to the model’s internals (e.g., gradients or confidence scores). These settings are idealized and do not correspond directly to practical attack conditions. However, these attacks are included for two important reasons. First, they allow us to assess the internal robustness of our model under strong adversarial conditions, thereby providing insight into its behavior in worst-case scenarios. For example, the gradient-based attack assumes white-box access and is designed to expose vulnerabilities in the learned feature space by directly manipulating model-sensitive directions. The genetic algorithm-based attack simulates a black-box adversary with query access to confidence scores, reflecting a more restricted but still potent threat model. Random noise attacks, while non-adaptive, serve as a baseline to evaluate the model’s resilience to generic feature perturbations. Second, such feature-space attacks are commonly adopted in the adversarial robustness studies for stress-testing models under controlled conditions. They are not intended to simulate realistic attacks, but rather to reveal how easily a classifier may be fooled when the attacker can exploit flaws in its learned decision boundaries.
@@
In addition, we include the TRA–$L_{0}$ distance curves for RTC and CFTC under GradArgmax and GeneticAlg attacks, as shown in Figure 13. Here, the $L_{0}$ distance represents the number of perturbed Hidost features required for a successful attack, serving as a measure of perturbation strength. This is conceptually similar to the Relative Perturbation Ratio (RPR) discussed in §$\S$6.5.
@@ I. Detailed Results of Ablation Study
The detailed results of the ablation study are presented in Table 12. Removing PDFObj IR and applying embeddings directly on raw content resulted in a decline in the classifier’s performance. Additionally, we selected models in §$\S$6.5 that exhibited a certain level of adversarial robustness: BERT-20k PV-DM-65k, BERT-65k, and text-embedding-3. Subsequently, we removed the ORG structure and used the average of node vectors for classification, and the classification results are shown in Table 12. The results indicate that after removing the ORG structure, the models that originally had adversarial capabilities no longer detected any adversarial samples, suggesting that the ORG structure contributes to improving the classifier’s adversarial robustness. Additionally, we observed that following the removal of the ORG structure, apart from BERT-65k, the combinations of other embedding schemes with DNN classifiers experienced a certain degree of decline in classification metrics compared to the combinations with GIN classifiers. This underscores that our trained BERT-65k scheme is more accurate in representing the semantics of PDF objects.

### v2

@@ 6.6. Ablation Study
Impact of Embedding Raw Content on Classification In this experiment, we first pre-trained the raw content of the PDF object, shown in Figure 1(b), using the standard BERT pre-training method with default preprocessing and tokenization configurations. This pre-training was performed on the baseline dataset (with a corpus size of 20k samples). Then, we applied the general embedding schemes from $\S$5.3 to embed the raw content of PDF objects, obtaining node embeddings. We then evaluated their performance on the PDF malware classification task, maintaining the same GIN model and training settings as described in $\S$5.4. The results, presented in the first three rows of Table 6, indicate that the BERTR scheme, pre-trained on raw content, performs worse than the BERT scheme pre-trained on PDFObj IR (as shown in Table 4). Specifically, with a corpus size of 20k, the accuracy on the baseline dataset decreased by 0.42%, and on the extended dataset, it dropped by 2.07%. Furthermore, the results show that directly applying general embedding schemes to raw content performs worse than applying them at the PDFObj IR level. For instance, the BERT BaseR scheme demonstrated a significant performance drop: compared to BERT Base in Table 4, accuracy on the baseline and extended datasets decreased by 1.29% and 8.56%, respectively. Similarly, CodeT5R saw accuracy reductions of 0.18% and 2.84%, respectively. Detailed results are provided in Appendix I.
@@
Impact of ORG Structure on Classification To evaluate the impact of graph structures on classifier robustness, we removed the ORG structure and GIN classifier. Given the variability in the number of objects across PDF samples, we represented each PDF by averaging the semantic vectors of all objects and used this representation as input to a DNN classifier. We selected four PDFObj2Vec schemes that demonstrated some adversarial robustness, as discussed in $\S$6.5. We trained the DNN classifier on the baseline dataset and assessed its performance. The experimental results highlight the critical importance of the ORG structure in enhancing both the performance and robustness of PDF malware classification. When the ORG structure was removed and a DNN classifier was used with semantic vectors of PDF objects, the classification accuracy on both the baseline and extended datasets declined. Furthermore, the DNN classifiers completely failed to exhibit resistance to adversarial samples, rendering them unable to detect such attacks effectively.
@@ 7. Discussion & Conclusion
Obfuscation Obfuscation is a common tactic in malware to evade detection, and similar techniques, such as keyword obfuscation, are used in PDF malware to bypass detection mechanisms (obs_pdfs_sentinelone; lu2013obfuscation; Pires_medium). For example, replacing Name /URI with /#55RI, prevents classifiers that rely on keyword statistical features from obtaining correct statistical features and also hinders analyzers that use Hidost path features from extracting the correct path features. Our method mitigates such obfuscation by retaining the obfuscated keyword data in the PDFObj IR corpus, allowing us to learn the obfuscated semantics. Additionally, we integrate the structural features of ORG to classify PDF malware, enhancing the precise detection of obfuscated PDF malware.
Concept Drift Over time, the effectiveness of traditional machine learning models trained on outdated datasets tends to decline (jordaney2017transcend; yang2021cade; barbero2022transcending). As multimedia evolves, benign PDFs are advancing faster than malicious ones, altering their characteristics significantly. Consequently, SOTA classifiers, including RTC (chen2020training) and CFTC (tong2019improving), struggle to accurately identify new benign samples of the extended dataset, leading to lower accuracy, as shown in Table 4. Since PDFObj2Vec is trained on PDFObj IR, a fundamental representation of PDF objects, the basic changes in PDFObj IR are minimal, regardless of variations in the PDF structure. Therefore, our proposed method is minimally affected by concept drift. The experimental results presented in Table 4 demonstrate that our proposed approaches, which leverage PDFObj IR, not only outperform methods that do not utilize this representation but also surpass SOTA classifiers across all metrics on the most recent extended dataset.
@@ B. Error Handling and Correction
Apart from the three main errors discussed in $\S$4.3, there are other errors that we have identified. These additional errors and their handling measures are listed in Table 8.
@@ C. Details of Pre-trained PDFObj2Vec
Word2Vec-Based Scheme For a sentence structure like Type_name Parent_ref MediaBox_num_list Resources_dict Resources/P-rocSet_name_list Resources/XObject_DICT... , suppose we select MediaBox_num_list as the target word, and the context window size is 2. This means we will use the two words before and after this word as the context, as shown in Figure 11. Thus, the context words are: [Type_name Parent_ref Resources_dict Resources/ProcSet_name_list]. The overall semantics of the context are represented by calculating the average of the context word vectors. This context vector is then fed into a softmax classifier, which predicts the probability of each word being the target word. In $\S$6, our implementation of Word2Vec consists of two embedding layers, one for learning the embeddings of the center words and the other for the context words, along with a Dropout layer. We update the word vectors using the cross-entropy loss between the target words and the model’s predictions. The optimizer used is Stochastic Gradient Descent (SGD), with a learning rate of 5e-4.
@@ E. Details of GIN Classifier
We trained our implemented GIN classifier on the baseline dataset using the Deep Graph Library (gdlai). The classifier consists of two GIN layers, each with a hidden layer size of 256, employing mean aggregation and the ReLU activation function. The update rule for a GIN graph convolution layer can be expressed as:
@@ H. Additional Details of Adversarial Attack
We summarize the adversarial attacks we use in $\S$6.5, as shown in Table 11. Specifically, the gradient-based attack, genetic algorithm-based attack, and random noise attack are conducted in the feature space, where attackers are assumed to have varying degrees of access to the model’s internals (e.g., gradients or confidence scores). These settings are idealized and do not correspond directly to practical attack conditions. However, these attacks are included for two important reasons. First, they allow us to assess the internal robustness of our model under strong adversarial conditions, thereby providing insight into its behavior in worst-case scenarios. For example, the gradient-based attack assumes white-box access and is designed to expose vulnerabilities in the learned feature space by directly manipulating model-sensitive directions. The genetic algorithm-based attack simulates a black-box adversary with query access to confidence scores, reflecting a more restricted but still potent threat model. Random noise attacks, while non-adaptive, serve as a baseline to evaluate the model’s resilience to generic feature perturbations. Second, such feature-space attacks are commonly adopted in the adversarial robustness studies for stress-testing models under controlled conditions. They are not intended to simulate realistic attacks, but rather to reveal how easily a classifier may be fooled when the attacker can exploit flaws in its learned decision boundaries.
@@
[…]
@@ I. Detailed Results of Ablation Study
The detailed results of the ablation study are presented in Table 12. Removing PDFObj IR and applying embeddings directly on raw content resulted in a decline in the classifier’s performance. Additionally, we selected models in $\S$6.5 that exhibited a certain level of adversarial robustness: BERT-20k PV-DM-65k, BERT-65k, and text-embedding-3. Subsequently, we removed the ORG structure and used the average of node vectors for classification, and the classification results are shown in Table 12. The results indicate that after removing the ORG structure, the models that originally had adversarial capabilities no longer detected any adversarial samples, suggesting that the ORG structure contributes to improving the classifier’s adversarial robustness. Additionally, we observed that following the removal of the ORG structure, apart from BERT-65k, the combinations of other embedding schemes with DNN classifiers experienced a certain degree of decline in classification metrics compared to the combinations with GIN classifiers. This underscores that our trained BERT-65k scheme is more accurate in representing the semantics of PDF objects.